# Improving the Cosmological Constraints by Inferring the Formation Channel of Extreme-mass-ratio Inspirals

### v1

@@ 4.2 Constraints on $w_{0}$ and $w_{a}$
The errors on the dark energy EoS parameters $w_{0}$ and $w_{a}$ are shown in Figure 4. Here we fix $h=0.678$ and $\Omega_{M}=0.307$. We find that only $w_{0}$ can be effectively constrained and $w_{a}$ is not constrained in almost all population models except in the M7 and M12 models. The much better constraint on $w_{0}$ than on $w_{a}$ implies that inferring the deviation of the dark energy EoS with respect to the value $w=-1$ of the standard model requires much better EMRI GW detections than reconstructing the evolution of the dark energy EoS with redshift. Even under the M7 and M12 models, despite the fact that they have optimistic EMRI rates such that TianQin I+II and TianQin+LISA can detect a total number of more than one thousand EMRIs, still, a stringent constraint on $w_{a}$ cannot be obtained.
Table I: Expected relative errors on the $\Lambda$CDM cosmological model parameters $(H_{0},\Omega_{M})$ and the CPL dark energy model parameters $(w_{0},w_{a})$ for the 11 EMRI population models and assuming three different detector configurations: TianQin, TianQin I+II and TianQin+LISA. The bar “--” means no effective constraint.
@@
More specifically, under the fiducial population model, the error of $w_{0}$ can reach about $37\%$ for TianQin, about $12\%$ for TianQin I+II, and about $6.5\%$ for TianQin+LISA. Under the other population models, i.e., from M2 to M12 except M11, the expected errors of $w_{0}$ can reach $6.7\%-48\%$ for TianQin, about $5.4\%-43\%$ for TianQin I+II, and about $4.2\%-22\%$ for TianQin+LISA. Similar to the constraints on $H_{0}$, the errors on $w_{0}$ also vary considerably across population models. Finally, we also summarize the expected constraint precisions on $w_{0}$ and $w_{a}$ for various cases in Table I.
5 Constraints on the formation channel of EMRIs
So far, we have assumed that each candidate host galaxy has the same probability of hosting an EMRI, whether it is a normal galaxy or an AGN. Now we drop the assumption and use the statistical framework described in Section 2.3 to constrain $f_{\rm agn}$, which is the fraction of EMRIs which reside in AGNs. To account for the random fluctuation of the AGN number density in the Universe, we use Monte-Carlo simulations to estimate the background distribution $P_{\rm bg}(\lambda)$ of the likelihood ratio $\lambda$ in Equation (18). We assume that $100$ EMRIs will be detected by the TianQin+LISA network (the number is expected to be reached with 5 years of detecting under most population models, except for the M5, M8, and M11 models, see Fan et al., 2020). For example, several background distributions for different $f_{\rm agn}$ are shown in Figure 5 as the dashed lines. In addition, to account for the randomness in simulated realizations, we also estimate the probability distribution $P(\lambda)$ through several times of independent simulations, as shown by the solid lines in Figure 5. We can see that the variation of $P(\lambda)$ is more significant than $P_{\rm bg}(\lambda)$ as $f_{\rm agn}$ increases, this is because the variation of $P(\lambda)$ is caused by a combination of changes in the value of $f_{\rm agn}$ and the numbers of candidate host AGNs $\{N_{{\rm agn},i}\}$ whereas the variation of $P_{\rm bg}(\lambda)$ is only caused by the change of $f_{\rm agn}$. Since the statistical significance mainly depends on the number of GW sources, we denote $N_{{\rm GW},3\sigma}^{\rm threshold}$ as the minimum number of detected EMRIs that is required to reject the null hypothesis at a significance level of $3\sigma$, and we study its dependence on the value of $f_{\rm agn}$.
Figure 5: Typical background distribution $P_{\rm bg}(\lambda)$ (dashed line) and probability distribution $P(\lambda)$ (solid line) of the likelihood ratio $\lambda$ in Equation (18) for 100 EMRIs detected by TianQin+LISA, under the M1 population model with fixed cosmology. The colors black, blue, cyan, and red represent four fractions of $f_{\rm agn}=0.1,0.3,0.6,1.0$, respectively.
@@ 5 Constraints on the formation channel of EMRIs
Figure 6: Minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ of detected EMRI events required to testify the EMRI-AGN correlation with a $3\sigma$ significance. Here we have used the M1 population model. The black, blue, and red error bars and lines correspond to TianQin, TianQin I+II, and TianQin+LISA observations. The solid lines represent the case that both $H_{0}$ and $\Omega_{M}$ are fix to the true values, and the dashed lines with corresponding error bars represent the case in which both $H_{0}$ and $\Omega_{M}$ are freely variable within the prior. The error bars correspond to $1\sigma$ CIs.
@@
Here, we take the M1 model as the fiducial population model again to estimate the required minimum number of detected EMRIs. When we convert the error of luminosity distance to the redshift error using Equations (19a) and (19b), we assume two cosmological conditions, namely: (i) “fixed cosmology” that the cosmological parameters are fixed to their true values, and (ii) “variable cosmology” that the cosmological parameters $H_{0}$ and $\Omega_{M}$ are freely variable within their priors. The latter condition is a conservative assumption, considering that there is a tension in current measurements of the Hubble constant (Freedman, 2017; Aghanim et al., 2020; Riess et al., 2021; Di Valentino et al., 2021; Perivolaropoulos & Skara, 2022). The resulting dependence of the required minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ on the fraction $f_{\rm agn}$ is shown in Figure 6. We can see that $N_{{\rm GW},3\sigma}^{\rm threshold}$ decreases with increasing $f_{\rm agn}$ in all cases.
@@
In the “fixed cosmology” condition, TianQin would need to detect at least $465$ EMRIs to be able to reject the null hypothesis and establish the EMRI-AGN correlation if $f_{\rm agn}=1$. Under the same condition, TianQin I+II needs to detect at least about $196$ EMRIs and TianQin+LISA needs to detect at least about $26$ EMRIs. In the “variable cosmology” condition, if $f_{\rm agn}=1$ the minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ required for TianQin, TianQin I+II, and TianQin+LISA are about $1650$, $584$, and $76$, respectively. The increase of $N_{{\rm GW},3\sigma}^{\rm threshold}$ in the case of variable cosmology is caused by an increase of the redshift range when transforming the luminosity distance range into the redshift space. For different cases, $N_{{\rm GW},3\sigma}^{\rm threshold}$ as a function of $f_{\rm agn}$ can be approximate as $N_{{\rm GW},3\sigma}^{\rm threshold}\propto f_{\rm agn}^{-1.2}$.
For the other EMRI population models, since the probability distributions of spatial localization errors of EMRIs detected by LISA and TianQin are not much different from the M1 model (Babak et al., 2017; Fan et al., 2020), the required minimum number of EMRI detections estimated in the M1 model also applies approximately to the other population models.
@@ 6 Discussions
6.1 Advantage of combining EMRI detections and AGN observations
Once either of the following two conditions is met, we can use AGN catalogs to extract the redshift information of the detected EMRIs (or at least assign higher weights to the AGNs): (i) the fraction of the EMRIs originate in AGNs, $f_{\rm agn}$, is determined to be close to $1$ by statistical analysis of the EMRI-AGN correlation; and (ii) a detected EMRI has a zero eccentricity at the plunge (see, e.g., Pan et al., 2021; Amaro-Seoane et al., 2023, for explanation). We note that LISA and TianQin can measure the eccentricities of EMRIs to a precision of $\Delta e\approx 10^{-6}$ (Babak et al., 2017; Fan et al., 2020), so they can differentiate the formation channels.
@@ 6.1 Advantage of combining EMRI detections and AGN observations
Figure 7: Constraints on $(h,\Omega_{M})$ (left panel) and $(w_{0},w_{a})$ (right panel) by combining EMRI detections of TianQin with AGN catalogs. The eccentricities at plunges are all set to zero, and the other parameters are derived from the M1 model. The blue contours represent the constraints derived from candidate host galaxies and the red contours represent the constraints from candidate host AGNs, respectively. The two levels of contour correspond to $1\sigma$ and $2\sigma$ CIs. The cyan crosses represent the true values of the cosmological parameters used in our simulations.
@@
In addition, we would like to point out that EMRI GW sources may be accompanied by observable EM signals (Wang et al., 2019b, 2022b; Li et al., 2023; Amaro-Seoane et al., 2023). As long as the EM signals from EMRI sources do exist, current and planning time-domain survey telescopes, such as the Wide Field Survey Telescope (WFST) (Wang et al., 2023; Huang et al., 2024), the Multi-channel Photometric Survey Telescope (Mephisto) (Yuan et al., 2020), Vera Rubin Observatory (also known as LSST; Ivezić et al., 2019), the Einstein Probe (Yuan et al., 2015), Athena (Nandra et al., 2013; McGee et al., 2020), the Five-Hundred Aperture Spherical Radio Telescope (FAST) (Nan et al., 2011), the Square Kilometre Array (SKA) (Janssen et al., 2015), and so on, could provide promising searches for the EM signals. Once the EM signals of EMRIs are observed, they could help us narrow down the pools of the candidate host galaies or even uniquely identify the host galaxies of the EMRIs. This will allow us to not only improve the precisions of the cosmological parameters by more accurately measuring the redshifts of the EMRIs but also enhance our understanding for the formation channel of the EMRIs.
@@ 6.2 Effects of eccentricity on the spatial localization of EMRIs
Since we have set the eccentricities of the EMRIs to be uniformly distributed in the range of 00 to $0.2$ when simulating the EMRI detections, however, the EMRIs originated in AGNs are expected to have a zero eccentricity at plunge (Pan et al., 2021; Amaro-Seoane et al., 2023). So that if there is a fraction of $f_{\rm agn}$ of EMRIs that do originate in AGNs, the values of eccentricities of EMRIs may affect the measurement precisions of their source parameters. Here, we analyze the effect of eccentricities of EMRIs at plunges on their spatial localization errors, the result is shown in Figure 8.
From Figure 8, we can see that the effect of EMRIs’ eccentricities at plunges on their spatial localization errors (sky position and distance) are not significant, and the error increases with the decrease of the eccentricity very slightly. This implies that both the expected errors on the cosmological parameters presented in Section 4 and the required minimum number of EMRI detections for establishing the EMRI-AGN correlation presented in Section 5 will not be significantly changed.
Figure 8: Variation of the spatial localization errors of TianQin EMRIs caused by their eccentricities at plunges. The gray (red) violinplots represent the probability distributions for the sky localizations ($D_{L}$ estimations) of EMRIs, and the gray (red) horizontal dashed line represents the median of the first violinplot.
@@ 6.3 Prospect of clarifying the Hubble tension
The cosmological standard model—flat-$\Lambda$CDM, is challenged by the Hubble tension (Freedman, 2017; Di Valentino et al., 2021; Perivolaropoulos & Skara, 2022). The measurement of $H_{0}$ obtained by the flat-$\Lambda$CDM model plus cosmic microwave background observations is $67.4\pm 0.5~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}$, with a relative error of about $0.7\%$ (Aghanim et al., 2020), and the $H_{0}$ measurement obtained by the cosmic distance ladder calibrations and type Ia supernova observations is $73.2\pm 1.3~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}$, with a relative error of about $1.8\%$ (Riess et al., 2021). So if we want to clarify the Hubble tension by a new independent measurement, the independently measured $H_{0}$ needs to be better than a precision of about $2\%$.
We can see from Table I that TianQin does not achieve a precision of 2% in constraining $H_{0}$ under all EMRI population models. As for TianQin I+II, it achieves a precision better than $2\%$ only under the population models with optimistic EMRI rates, such as the M7 and M12 models. However, the multi-detector network formed by LISA and TianQin, i.e., TianQin+LISA, can much better constrain $H_{0}$. For example, under most of the population models, except for the M8 and M11 models, the TianQin+LISA network can constrains $H_{0}$ to a precision of near or better than $2\%$. Therefore, the results of our work support a joint observation of GWs by TianQin and LISA.

### v2

@@ 4.2 Constraints on $w_{0}$ and $w_{a}$
The errors on the dark energy EoS parameters $w_{0}$ and $w_{a}$ are shown in Figure 5. Here we fix $h=0.678$ and $\Omega_{M}=0.307$. We find that only $w_{0}$ can be effectively constrained and $w_{a}$ is not constrained in almost all population models. Even under the M7 and M12 models, despite the fact that they have optimistic EMRI rates such that TianQin I+II and TianQin+LISA can detect a total number of more than one thousand EMRIs, still, a stringent constraint on $w_{a}$ cannot be obtained. The much better constraint on $w_{0}$ than on $w_{a}$ implies that reconstructing the evolution of the dark energy with redshift requires much better EMRIs than those needed to infer the deviation of the dark-energy EoS with respect to the value $w=-1$ of the standard model.
More specifically, under the fiducial population model, the error of $w_{0}$ can reach about $37\%$ for TianQin, about $12\%$ for TianQin I+II, and about $6.5\%$ for TianQin+LISA. Under the other population models, i.e., from M2 to M12 except M11, the expected errors of $w_{0}$ can reach $6.7\%-48\%$ for TianQin, about $5.4\%-43\%$ for TianQin I+II, and about $4.2\%-22\%$ for TianQin+LISA. Similar to the constraints on $H_{0}$, the errors on $w_{0}$ also vary considerably across population models. Finally, in Table I we also summarize the expected precisions in the constraints on $w_{0}$ and $w_{a}$ for various models.
Table I: Expected relative errors of the $\Lambda$CDM cosmological model parameters $(H_{0},\Omega_{M})$ and the CPL dark energy model parameters $(w_{0},w_{a})$ for the 11 EMRI population models and assuming three different detector configurations: TianQin, TianQin I+II and TianQin+LISA. The bar “--” means no effective constraint.
@@
5 Constraining the formation channel of EMRIs
@@
Figure 6: Minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ of detected EMRI events required to testify the EMRI-AGN correlation with a $3\sigma$ significance. The black, blue, and red error bars and lines correspond to TianQin, TianQin I+II, and TianQin+LISA observations. The solid lines represent the case in which both $H_{0}$ and $\Omega_{M}$ are fix to the true values, and the dashed lines with corresponding error bars represent the case in which both $H_{0}$ and $\Omega_{M}$ are freely variable within the prior. The error bars correspond to $1\sigma$ CI.
So far, we have assumed that each candidate host galaxy has the same probability of hosting an EMRI, be it a normal galaxy or an AGN. Now we drop this assumption and use the statistical framework described in Section 2.3 to constrain $f_{\rm agn}$, which is the fraction of EMRIs which reside in AGNs. Since the statistical significance increases with the number of GW sources, we denote $N_{{\rm GW},3\sigma}^{\rm threshold}$ as the minimum number of detected EMRIs that is required to reject the null hypothesis at a significance level of $3\sigma$, and we study its dependence on the value of $f_{\rm agn}$.
@@
Here, we follow the works of Babak et al. (2017) and Pan et al. (2021) to generate the MBH mass function and its evolution with redshift. Correspondingly, the dimensionless MBH spin is set to $0.98$ and the mass of the stellar BH is fixed at $10~{}\!M_{\odot}$. The above choices are commonly used in the previous works (Barausse, 2012; Amaro-Seoane et al., 2013; Abbott et al., 2019b, 2023b), thus we adopt them for easier comparison with the previous works. The eccentricities of EMRIs are expected to be significantly different in different models. For the EMRIs in normal galaxies, we assumed a uniform distribution of the eccentricities at plunge within the range of $[0,0.2]$, as is proposed by Babak et al. (2017); for the EMRIs originating in AGNs, we set their eccentricities to zero, to account for the damping of eccentricity suggested by Pan et al. (2021). When we convert the error of luminosity distance to the redshift error using Equations (20a) and (20b), we assume two cosmological conditions, namely: (i) “fixed cosmology” in which the cosmological parameters are fixed to their true values, and (ii) “variable cosmology” where the cosmological parameters $H_{0}$ and $\Omega_{M}$ are freely variable within their priors. We consider the latter condition to account for the tension in the current measurements of the Hubble constant (Freedman, 2017; Aghanim et al., 2020; Riess et al., 2021; Di Valentino et al., 2021; Perivolaropoulos & Skara, 2022). The resulting dependence of the required minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ on the fraction $f_{\rm agn}$ is shown in Figure 6. We can see that $N_{{\rm GW},3\sigma}^{\rm threshold}$ decreases with increasing $f_{\rm agn}$ in all cases.
@@
In the “fixed cosmology” condition, TianQin would need to detect at least $1300$ EMRIs to be able to reject the null hypothesis and establish the EMRI-AGN correlation if $f_{\rm agn}=1$. Under the same condition, TianQin I+II needs to detect at least about $450$ EMRIs and TianQin+LISA needs to detect at least about $30$ EMRIs. In the “variable cosmology” condition, if $f_{\rm agn}=1$ the minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ required for TianQin, TianQin I+II, and TianQin+LISA are about $5000$, $1300$, and $90$, respectively. The increase of $N_{{\rm GW},3\sigma}^{\rm threshold}$ in the case of variable cosmology is caused by an increase of the redshift range when transforming the luminosity distance range into the redshift space. For different cases, $N_{{\rm GW},3\sigma}^{\rm threshold}$ as a function of $f_{\rm agn}$ can be approximate as $N_{{\rm GW},3\sigma}^{\rm threshold}\propto f_{\rm agn}^{-1.6}$.
Figure 7: The error in constraining $f_{\rm agn}$ as a function of the number of detected EMRIs. Here we fix the cosmological parameters to their true values. The black, blue, and red error bars correspond to TianQin, TianQin I+II, and TianQin+LISA. The thick and thin error bars correspond to $1\sigma$ and $2\sigma$ CIs, respectively. The dots in the middle of the error bars represent the expectations. The green dashed horizontal line represents the injected $f_{\rm agn}$ in the simulations.
After testing the EMRI-AGN correlation, we proceed to estimate the value of $f_{\rm agn}$ using Equation (19) and Monte-Carlo simulations. As an illustration, we set the fraction of EMRIs originating in AGNs to $f_{\rm agn}=0.7$. The variation of the error of $f_{\rm agn}$ with the number of detected EMRIs for different detector configurations is shown in Figure 7. We find that as the number of detected EMRIs approaches or exceeds the threshold number $N_{{\rm GW},3\sigma}^{\rm threshold}$ (refer to Figure 6), the constraint on $f_{\rm agn}$ converges toward the injected value, with the constraining error diminishing as the number of detected EMRIs increases. When the number of detected EMRIs is considerably lower than $N_{{\rm GW},3\sigma}^{\rm threshold}$, the constraint on $f_{\rm agn}$ is consistent with $f_{\rm agn}=0.01$ (the fraction of AGNs among all galaxies we assumed), but with large error. Furthermore, for an equivalent number of EMRI detections, the TianQin+LISA network significantly enhances the precision of constraining $f_{\rm agn}$. For instance, with 1000 EMRI detections, the errors corresponding to $1\sigma$ CIs are $\Delta f_{\rm agn}\sim 0.33$ for TianQin and $\Delta f_{\rm agn}\sim 0.21$ for TianQin I+II, whereas the error for TianQin+LISA shrinks to $\Delta f_{\rm agn}\sim 0.09$.
We would like to note that the real EMRI population in the universe is more completed than the modeled ones presented in this work. For examples, both the spins of the MBHs (Sesana et al., 2014; Bustamante & Springel, 2019; Sisk-Reynés et al., 2022) and the masses of the stellar-mass BHss (Pan et al., 2022; Abbott et al., 2023b) should follow certain distributions, but we have adopted fixed values for them. In addition, the EMRIs originating in normal galaxies (Hopman et al., 2007; Yunes et al., 2008; Berry & Gair, 2013; Amaro-Seoane, 2019; Han et al., 2020; Fan et al., 2022) or in those post-merger galaxies hosting MBHBs (Mazzolari et al., 2022; Naoz et al., 2022; Naoz & Haiman, 2023) may exhibit eccentricities much higher than $0.2$, and in AGNs the stellar-mass BHs would interact with the accretion disks (Yunes et al., 2011; Chatterjee et al., 2023). All these factors could affect the duration, GW amplitude, or extra modes of an EMRI, and consequently affect its detectability. Nevertheless, our work serves as a key step towards testing the EMRI-AGN correlation.
@@ 6 Discussions
6.1 Advantage of combining EMRIs with AGN observation
Once either of the following two conditions is met, we can use AGN catalogs to extract the redshift information of a detected EMRI (or at least assign higher weights to the AGNs): (i) the fraction of the EMRIs originate in AGNs, $f_{\rm agn}$, is statistically constrained to be close to $1$ by our method; and (ii) the detected EMRI has a zero eccentricity at the time of plunge (see, e.g., Pan & Yang, 2021; Pan et al., 2021; Amaro-Seoane et al., 2023, for explanation). We note that LISA and TianQin can measure the eccentricities of EMRIs to a precision of $\Delta e\approx 10^{-6}$ (Babak et al., 2017; Fan et al., 2020), so they can differentiate the formation channels.
@@ 6.1 Advantage of combining EMRIs with AGN observation
Figure 8: Constraints on $(h,\Omega_{M})$ (left panel) and $(w_{0},w_{a})$ (right panel) by combining the EMRIs detected by TianQin with AGN catalogs. The eccentricities at plunges are set to zero, and the other parameters are derived from the M1 model. The blue contours represent the constraints derived from candidate host galaxies and the red contours represent the constraints from candidate host AGNs. The two levels of contours correspond to $1\sigma$ and $2\sigma$ CIs, respectively. The cyan crosses mark the injected cosmological parameters in our simulations.
@@
In addition, we would like to point out that EMRIs may be accompanied by EM counterparts (Wang et al., 2019b, 2022b; Li et al., 2023; Amaro-Seoane et al., 2023). The EM signals could be detected by the current and planned time-domain survey telescopes, such as the Wide Field Survey Telescope (WFST) (Wang et al., 2023; Huang et al., 2024), the Multi-channel Photometric Survey Telescope (Mephisto) (Yuan et al., 2020), Vera Rubin Observatory (also known as LSST; Ivezić et al., 2019), the Einstein Probe (Yuan et al., 2015), Athena (Nandra et al., 2013; McGee et al., 2020), the Five-Hundred Aperture Spherical Radio Telescope (FAST) (Nan et al., 2011), or the Square Kilometre Array (SKA) (Janssen et al., 2015). Detecting an EM counterpart could help us narrow down the candidates host galaxies, or even uniquely identify the host galaxy of an EMRI. Such detections will enable us to not only improve the precisions of the cosmological parameters due to the more accurate redshift information of the EMRIs, but also better understand the formation channel of EMRIs.
@@ 6.2 Effects of eccentricity on the spatial localization of EMRIs
In our simulations, we have set the eccentricities of EMRIs (at plunge) to be uniformly distributed in the range of 00 to $0.2$ when simulating the EMRIs from normal galaxies, and set them to zero when the EMRIs come from AGNs (Pan & Yang, 2021; Pan et al., 2022, 2021; Amaro-Seoane et al., 2023). To evaluate whether the eccentricities of EMRIs may affect the precision in the measurement of their parameters, here we analyze the spatial localization errors for the EMRIs formed in different channels. The results are shown in Figure 9.
Figure 9: Spatial localization errors of the TianQin EMRIs as a function of their eccentricities at the time of plunge. The gray (red) violinplots represent the probability distributions for the sky localizations errors ($D_{L}$ estimation errors) of EMRIs, and the gray (red) horizontal dashed line shows the median of the corresponding error.
From Figure 9, we can see that the effect of eccentricity at plunge is weak on the spatial localization errors (sky position and distance). The weak dependence implies that neither the errors on the cosmological parameters presented in Section 4 nor the required minimum number of EMRIs for establishing the EMRI-AGN correlation presented in Section 5 is sensitive to our assumption of the distribution of EMRI eccentricities.
Although we have not considered in detail in this work, a special case worth mentioning is an EMRI whose eccentricity is close to unity. Such a system may occasionally form in a normal galaxy, and it would radiate multiple GW bursts before transitioning into the inspiraling phase (Rubbo et al., 2006; Hopman et al., 2007; Yunes et al., 2008; Berry & Gair, 2013; Amaro-Seoane, 2019; Han et al., 2020; Fan et al., 2022). The SNR of this extreme-mass-ration burst (EMRB) may exceed the detection threshold, allowing identification of individual bursts (Rubbo et al., 2006; Yunes et al., 2008; Han et al., 2020; Fan et al., 2022; Wu et al., 2023). Detecting these burst prior to an EMRI signal can enhance the spatial localization of the EMRI by partially breaking the degeneracy between parameters (Barack & Cutler, 2004; Mikóczi et al., 2012; Xuan et al., 2023). This prospect deserves future investigation.
@@ 6.3 Prospect of clarifying the Hubble tension
Recently, the cosmological standard model—flat-$\Lambda$CDM, is challenged by the Hubble tension (Freedman, 2017; Di Valentino et al., 2021; Perivolaropoulos & Skara, 2022). The $H_{0}$ obtained from the flat-$\Lambda$CDM model plus cosmic microwave background is $67.4\pm 0.5~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}$, with a relative error of about $0.7\%$ (Aghanim et al., 2020), and the $H_{0}$ from the cosmic distance ladders and type Ia supernovae is $73.2\pm 1.3~{}\!{\rm km}~{}\!{\rm s}^{-1}~{}\!\!{\rm Mpc}^{-1}$, with a relative error of about $1.8\%$ (Riess et al., 2021). So if we want to clarify the Hubble tension by a new independent measurement, the independently measured $H_{0}$ needs to be better than a precision of about $2\%$.
We can see from Table I that TianQin does not achieve a precision of 2% in constraining $H_{0}$ under the EMRI population models considered in this paper. As for TianQin I+II, it achieves a precision better than $2\%$ only under the population models with optimistic EMRI rates, such as the M7 and M12 models. However, the multi-detector network formed by LISA and TianQin, i.e., TianQin+LISA, can much better constrain $H_{0}$. For example, under most of the population models, except for the M8 and M11 models, the TianQin+LISA network can constrains $H_{0}$ to a precision of near or better than $2\%$. Therefore, our results support a joint observation of GWs by TianQin and LISA.